An electron has a speed of 1.4E7 m/s. If an uncertainty 65 nm in its potition is acceptable, what is the minimum possible percentage uncertianty in its linear momentum?
Answer #1

V = 1.4*107 AX*AP -h /2 =h /2 *1/AX ΔΡ AX = 65 nm 65*10^-9 m so h /2 1/AX AP= 6.626*10^34/ (4*pi) * 1/( 65*10^-9) = 8.11201x